# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### v2

@@ D.1 Setup Details
For simplicity, we assume that the nearest neighbor network is known and set the relevant hyperparameters accordingly. We took care to set $\gamma$ such that our method performs a similar amount of shrinkage compared to other methods, such as IntelPooling, which effectively uses a separate penalty matrix for users and time. To do so, we set a separate value of $\gamma$ for both users and time and set it to the maximum eigenvalue of the penalty matrix (random effect precision matrix) used by IntelPooling. We use 5 neighbors within the DML methods and set the other hyperparameters as follows: $\sigma=1$, $\lambda=1$, and $\delta=0.01$.
For the Neural-Linear method, we generate a $125\times 125$ array of baseline rewards (no action-specific component) to train the neural network prior to running the bandit algorithm. Consequently, the results shown in the paper for Neural-Linear are better than would be observed in practice because we allowed the Neural-Linear method to leverage data that we did not make available to the other methods. This setup offers the computational benefit of not needing to update the neural network within bandit replications, which substantially reduces the necessary computation time.
@@
Our neural networks consisted of four hidden layers with 10, 20, 20, and 10 nodes, respectively. The first two employ the ReLU activation function (Nair & Hinton, 2010) while the latter two employ the hyperbolic tangent. We chose to use the hyperbolic tangent for the last two layers because Snoek et al. (2015) found that smooth activation functions such as the hyperbolic tangent were advantageous in their neural bandit algorithm. The loss function was the mean squared error between the neural networks’ output and the baseline reward on a simulated data set. We trained our networks using the Adam optimizer (Kingma & Ba, 2014) with batch sizes of 200 for between 20 and 50 epochs. We simulated a separate validation data set to ensure to check that our model had converged and was generating accurate predictions.
Figure 5 compares the true baseline reward function in the nonlinear setting (left) to that estimated by the neural network (right). We see that neural network produced an accurate approximation of the baseline reward, which helps explain the good performance of the Neural-Linear method relative to other baseline approaches, such as Standard.
@@
D.2 Pairwise Comparisons for Main Simulation
@@
The DML-TS-NNR-BLM and DML-TS-NNR-BT methods perform well across all three settings and the difference between them is statistically indistinguishable. These methods perform about equally well compared to IntelPooling in the first setting, but they perform better in the other settings because the DML methods (1) can leverage network information and (2) can accurately model the nonlinear baseline reward in the third setting.
DML-TS-NNR-BLM and DML-TS-NNR-BT perform especially well in the nonlinear setting, the more general setting for which they were designed. They achieve pairwise win percentages of 91% and 87%, respectively, compared to only 48% for the next-best non-DML method (IntelPooling).
@@
1 2 3 4 5 6 7 Avg
1. IntelPooling - 54% 46% 58% 100%* 100%* 100%* 76%
2. DML-TS-NNR-BLM 46% - 50% 52% 100%* 100%* 100%* 75%
3. DML-TS-NNR-BT 54% 50% - 58% 100%* 100%* 100%* 77%
4. DML-TS-SU-BT 42% 48% 42% - 100%* 100%* 100%* 72%
5. AC 0%* 0%* 0%* 0%* - 0%* 0%* 0%
6. Standard 0%* 0%* 0%* 0%* 100%* - 0%* 17%
7. Neural-Linear 0%* 0%* 0%* 0%* 100%* 100%* - 33%
@@
1 2 3 4 5 6 7 Avg
1. IntelPooling - 0%* 10%* 10%* 100%* 100%* 84%* 51%
2. DML-TS-NNR-BLM 100%* - 54% 66%* 100%* 100%* 100%* 87%
3. DML-TS-NNR-BT 90%* 46% - 64% 100%* 100%* 100%* 83%
4. DML-TS-SU-BT 90%* 34%* 36% - 100%* 100%* 100%* 77%
5. AC 0%* 0%* 0%* 0%* - 0%* 0%* 0%
6. Standard 0%* 0%* 0%* 0%* 100%* - 0%* 17%
7. Neural-Linear 16%* 0%* 0%* 0%* 100%* 100%* - 36%
@@
1 2 3 4 5 6 7 Avg
1. IntelPooling - 2%* 6%* 24%* 100%* 94%* 62%* 48%
2. DML-TS-NNR-BLM 98%* - 56% 94%* 100%* 100%* 100%* 91%
3. DML-TS-NNR-BT 94%* 44% - 84%* 100%* 100%* 98%* 87%
4. DML-TS-SU-BT 76%* 6%* 16%* - 100%* 100%* 90%* 65%
5. AC 0%* 0%* 0%* 0%* - 0%* 0%* 0%
6. Standard 6%* 0%* 0%* 0%* 100%* - 10%* 19%
7. Neural-Linear 38%* 0%* 2%* 10%* 100%* 90%* - 40%
Table 1: Pairwise comparisons between methods in the three settings of the main simulation. Each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The full DML methods (DML-TS-NNR-BLM and DML-TS-NNR-BT) perform best in all three settings and their final regret is statistically indistinguishable.
D.3 Simulation with Rectangular Data Array
The main simulation involves simulating data from a triangular data array. At the 125-th (final) stage, the algorithm has observed 125 rewards for user 1, 124 rewards for user 2, and so on.
@@ D.3 Simulation with Rectangular Data Array
Figure 6: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. Similar to Figure 1, the full DML methods (DML-TS-NNR-BLM and DML-TS-NNR-BT) are highly competitive in all three settings and substantially outperform the other methods in the nonlinear setting.
The cumulative regret for these methods as a function of time—not stage—is shown in Figure 6. The qualitative results are quite similar to those from the main simulation. DML-TS-NNR-BLM and DML-TS-NNR-BT are highly competitive in all three scenarios and substantially outperform the other methods in the nonlinear setting; in fact, compared to the main simulation, the improvement of these methods over the others is even larger.
@@
1 2 3 4 5 6 7 Avg
1. IntelPooling - 62% 60% 58% 100%* 100%* 100%* 80%
2. DML-TS-NNR-BLM 38% - 50% 44% 100%* 100%* 100%* 72%
3. DML-TS-NNR-BT 40% 50% - 44% 100%* 100%* 100%* 72%
4. DML-TS-SU-BT 42% 56% 56% - 100%* 100%* 100%* 76%
5. AC 0%* 0%* 0%* 0%* - 0%* 0%* 0%
6. Standard 0%* 0%* 0%* 0%* 100%* - 0%* 17%
7. Neural-Linear 0%* 0%* 0%* 0%* 100%* 100%* - 33%
@@
1 2 3 4 5 6 7 Avg
1. IntelPooling - 0%* 0%* 8%* 100%* 100%* 86%* 49%
2. DML-TS-NNR-BLM 100%* - 48% 100%* 100%* 100%* 100%* 91%
3. DML-TS-NNR-BT 100%* 52% - 100%* 100%* 100%* 100%* 92%
4. DML-TS-SU-BT 92%* 0%* 0%* - 100%* 100%* 100%* 65%
5. AC 0%* 0%* 0%* 0%* - 0%* 0%* 0%
6. Standard 0%* 0%* 0%* 0%* 100%* - 0%* 17%
7. Neural-Linear 14%* 0%* 0%* 0%* 100%* 100%* - 36%
@@
1 2 3 4 5 6 7 Avg
1. IntelPooling - 0%* 0%* 12%* 100%* 100%* 18%* 38%
2. DML-TS-NNR-BLM 100%* - 44% 100%* 100%* 100%* 100%* 91%
3. DML-TS-NNR-BT 100%* 56% - 100%* 100%* 100%* 100%* 93%
4. DML-TS-SU-BT 88%* 0%* 0%* - 100%* 100%* 64% 59%
5. AC 0%* 0%* 0%* 0%* - 0%* 0%* 0%
6. Standard 0%* 0%* 0%* 0%* 100%* - 0%* 17%
7. Neural-Linear 82%* 0%* 0%* 36% 100%* 100%* - 53%
Table 2: Pairwise comparisons between methods in the three settings of the simulation with a rectangular array of data. As in Table 1, each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The full DML methods (DML-TS-NNR-BLM and DML-TS-NNR-BT) perform best in all three settings in terms of pairwise win percentages and are not significantly different from each other.
Table 2 displays the results from pairwise method comparisons, similar to those shown in Table 1. Again the results are qualitatively similar to those from the main simulation. In 26/30 pairwise comparisons, DML-TS-NNR-BLM and DML-TS-NNR-BT outperform the other methods in 100% of replications. The four remaining pairwise comparisons are not statistically significant and result from comparisons with DML-TS-SU-BT and IntelPooling in the “Homogeneous Users” setting. DML-TS-NNR-BLM and DML-TS-NNR-BT offer little or no benefit compared to these methods in that setting because (1) the baseline is linear, (2) no network information is available, (3) no time effects are present, and (4) partial pooling (employed by DML-TS-NNR-BLM, DML-TS-NNR-BT, and IntelPooling) offers no benefit compared to full pooling (employed by DML-TS-SU-BT) because the causal effects are exactly the same among users.
In summary, our methods substantially outperform the other methods in complex, general settings and perform competitively with the other methods in simple settings.
Appendix E Additional Details for Valentine Study
Personalizing treatment delivery in mobile health is a common application for online learning algorithms. We focus here on the Valentine study, a prospective, randomized-controlled, remotely administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients (Jeganathan et al., 2022). We aim to use smart watch data (Apple Watch and Fitbit) obtained from the Valentine study to learn the optimal timing of notification delivery given the users’ current context.
E.1 Data from the Valentine Study
Prior to the start of the trial, baseline data was collected on each of the participants (e.g., age, gender, baseline activity level, and health information). During the study, participants are randomized to either receive a notification ($A_{t}=1$) or not ($A_{t}=0$) at each of 4 daily time points (morning, lunchtime, mid-afternoon, evening), with probability 0.25. Contextual information was collected frequently (e.g., number of messages sent in prior week, step count variability in prior week, and pre-decision point step-counts).

### v3

@@ B.1 Setup Details
For simplicity, we assume that the nearest-neighbor network is known and set the relevant hyperparameters accordingly. We took care to set the hyperparameters such that DML-TS-NNR performs a similar amount of shrinkage compared to other methods, especially IntelPooling, which effectively uses a separate penalty matrix for users and time. To accomplish this, we used the same penalty matrices for DML-TS-NNR and IntelPooling, effectively generalizing the scalar $\gamma$ penalty discussed in the main paper. For the purposes of calculating the constant $B$ given in Algorithm 2, we used a separate value of $\gamma$ for both users and time and set them equal to the maximum eigenvalue of the corresponding penalty matrix. We use 5 neighbors within the DML methods and set the other hyperparameters as follows: $\lambda=1$, and $\delta=0.01$, and $v_{k}=1$ (i.e., $v_{k}$ is constant).
For the Neural-Linear method, we generate a $200\times 200$ array of baseline rewards to train the neural network prior to running the bandit algorithm. Consequently, the results shown in the paper for Neural-Linear are better than would be observed in practice because we allowed the Neural-Linear method to leverage data that we did not make available to the other methods. This setup offers the computational benefit of not needing to update the neural network within bandit replications, which substantially reduces the necessary computation time.
@@
Our neural networks consisted of four hidden layers with 10, 20, 20, and 10 nodes, respectively. The first two employ the ReLU activation function (Nair & Hinton, 2010) while the latter two employ the hyperbolic tangent. We chose to use the hyperbolic tangent for the last two layers because Snoek et al. (2015) found that smooth activation functions such as the hyperbolic tangent were advantageous in their neural bandit algorithm. The loss function was the mean squared error between the neural networks’ output and the baseline reward on a simulated data set. We trained our networks using the Adam optimizer (Kingma & Ba, 2014) with batch sizes of 200 for between 20 and 50 epochs. We simulated a separate validation data set to check that our model had converged and was generating accurate predictions.
Figure 5 compares the true baseline reward function in the nonlinear setting (left) to that estimated by the neural network (right). We see that the neural network produced an accurate approximation of the baseline reward, which helps explain the good performance of the Neural-Linear method relative to other baseline approaches, such as Standard.
@@
B.2 Additional Method Comparisons
Our results in the main paper indicate that DML-TS-NNR outperforms existing methods, especially in the complex longitudinal setting that we are targeting. These positive results beg the question: What aspects of DML-TS-NNR contribute most to its superior performance?
In this section, we describe an additional simulation study designed to answer this question. In it, we compare DML-TS-NNR (as implemented in the main paper) to three additional comparison methods: DML-TS-NNR-BLM, DML-TS-SU, and NNR-Linear. Each comparison isolates particular aspects of our algorithm to understand its contribution to DML-TS-NNR’s performance.
DML-TS-NNR-BLM is a version of DML-TS-NNR that employs Bagged Linear Models as the baseline supervised learning algorithm. Unlike the method implemented in the main paper (which uses bagged stochastic gradient trees), this baseline model is misspecified in the nonlinear setting; i.e., it cannot accurately model the nonlinear baseline reward function. Consequently, this comparison will allow us to gauge the impact of the supervised learning method used in our method. In particular, it should allow us to validate whether the algorithm is robust to baseline misspecification as the theory suggests.
DML-TS-SU is a Single-User version of our algorithm. Rather than having a separate $\theta_{i}$ for each user, this method estimates a single advantage function that is shared across all users. It does, however, employ DML, time effects, and nearest-neighbor regularization (for the time effects). Consequently, this comparison will help us understand the impact of the user-specific parameters in our simulation.
NNR-Linear is similar to the existing network-cohesion bandit algorithms described in the main paper. It includes a distinct $\theta_{i}$ for each user and regularizes them toward each other using a Laplacian penalty. It does not, however, use DML or time effects. As a result, this comparison isolates the impact of these two factors and should provide evidence that DML-TS-NNR is more broadly applicable than existing network-cohesion bandit algorithms.
Figure 6 displays the cumulative regret as a function of the stage for these methods across the three settings used in the main paper. We see that all methods perform similarly in panel (a), the setting with homogeneous users. However, in panel (b) we see that the regret achieved by DML-TS-SU begins to resemble a straight upward-sloping line because it cannot appropriately model the user-specific advantage functions.
In panel (c), we observe a much wider spread of performance, with DML-TS-NNR and DML-TS-NNR-BLM performing similarly and substantially outperforming the other methods. The fact that DML-TS-NNR and DML-TS-NNR-BLM perform similarly indicates that our algorithm is, in fact, robust to misspecification as expected. The comparison to NNR-Linear reveals that the DML and time effects—the new components of our method compared to existing network-cohesion bandit algorithms—do contribute meaningfully to DML-TS-NNR’s superior performance. We found in additional simulations studies not reported here that the gap between NNR-Linear and our method widens as the magnitude of the time effects increases.
Figure 6: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings for the additional comparison methods. The fact that DML-TS-NNR outperforms DML-TS-SU demonstrates the importance of the user-specific parameters. The fact that DML-TS-NNR outperforms NNR-Linear shows that previous network cohesion approaches do not adequately address the (1) nonlinear nature of the baseline reward and (2) presence of time effects. The comparable performance of DML-TS-NNR-BLM to DML-TS-NNR highlights how our algorithm is robust to misspecification of the baseline reward model.
B.3 Pairwise Comparisons
@@
DML-TS-NNR and DML-TS-NNR-BLM perform well across all three settings and the difference between them is statistically indistinguishable in 2 of 3 settings. These methods perform about equally well compared to IntelPooling and NNR-Linear in the first setting, but they perform better in the other two. In the Nonlinear setting in particular, DML-TS-NNR and DML-TS-NNR-BLM substantially outperform all other methods, and the pairwise differences are statistically significant at the 5% level.
@@
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 58%* 48% 52% 32%* 100%* 100%* 100%* 70%
2. DML-TS-NNR-BLM 42%* - 40%* 36%* 22%* 100%* 100%* 100%* 63%
3. DML-TS-SU 52% 60%* - 48% 34%* 100%* 100%* 100%* 71%
4. NNR-Linear 48% 64%* 52% - 32%* 100%* 100%* 100%* 71%
5. IntelPooling 68%* 78%* 66%* 68%* - 100%* 100%* 100%* 83%
6. Neural-Linear 0%* 0%* 0%* 0%* 0%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
@@
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 46% 64% 58% 90%* 100%* 100%* 100%* 80%
2. DML-TS-NNR-BLM 54% - 64%* 60% 100%* 100%* 100%* 100%* 83%
3. DML-TS-SU 36% 36%* - 44% 88%* 100%* 100%* 100%* 72%
4. NNR-Linear 42% 40% 56% - 94%* 100%* 100%* 100%* 76%
5. IntelPooling 10%* 0%* 12%* 6%* - 80%* 100%* 100%* 44%
6. Neural-Linear 0%* 0%* 0%* 0%* 20%* - 100%* 100%* 31%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
@@
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 54% 100%* 98%* 98%* 100%* 100%* 100%* 93%
2. DML-TS-NNR-BLM 46% - 94%* 98%* 100%* 100%* 100%* 100%* 91%
3. DML-TS-SU 0%* 6%* - 42% 78%* 100%* 100%* 100%* 61%
4. NNR-Linear 2%* 2%* 58% - 82%* 100%* 100%* 100%* 63%
5. IntelPooling 2%* 0%* 22%* 18%* - 98%* 100%* 100%* 49%
6. Neural-Linear 0%* 0%* 0%* 0%* 2%* - 98%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 2%* - 100%* 15%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Table 2: Pairwise comparisons between methods in the three settings of the main simulation. Each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. DML-TS-NNR and DML-TS-NNR-BLM perform well in all three settings, and their final regret is statistically indistinguishable in 2 of 3 settings. They substantially outperform all other methods in the Nonlinear setting.
B.4 Simulation with Rectangular Data Array
The main simulation involves simulating data from a triangular data array. At the 200-th (final) stage, the algorithm has observed 200 rewards for user 1, 199 rewards for user 2, and so on.
@@ B.4 Simulation with Rectangular Data Array
Figure 7: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. Similar to Figure 1, DML-TS-NNR is competitive in the first setting and substantially outperforms the competitors in the other settings.
Figure 8: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. We observe the same performance ordering as in Figure 6, but here the relative differences are even larger.
The cumulative regret for these methods as a function of time—not stage—is shown in Figure 7. Qualitatively, the results are nearly identical to those from the main simulation (compare to Figure 1). The differences in final regret are even larger in this simulation than they were in the main simulation, presumably because this study exhibits greater variation in time effects due to the balance across time.
Figure 8 displays comparisons similar to those shown in Figure 6 but for the rectangular simulation. We again see that the results are qualitatively similar to the main simulation. The methods perform similarly in the Homogeneous Users setting. DML-TS-SU performs poorly in the Heterogeneous Users setting, but the other methods perform similarly. Finally, DML-TS-NNR and DML-TS-NNR-BLM substantially outperform the other methods in the nonlinear setting. Again, the differences in performance are even larger in this simulation than in the main simulation.
@@
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 54% 46% 42% 40% 100%* 100%* 100%* 69%
2. DML-TS-NNR-BLM 46% - 46% 38%* 32%* 100%* 100%* 100%* 66%
3. DML-TS-SU 54% 54% - 44% 34% 100%* 100%* 100%* 69%
4. NNR-Linear 58% 62%* 56% - 52% 100%* 100%* 100%* 75%
5. IntelPooling 60% 68%* 66% 48% - 100%* 100%* 100%* 77%
6. Neural-Linear 0%* 0%* 0%* 0%* 0%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
@@
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 56% 100%* 44% 100%* 100%* 100%* 100%* 86%
2. DML-TS-NNR-BLM 44% - 100%* 38% 100%* 100%* 100%* 100%* 83%
3. DML-TS-SU 0%* 0%* - 2%* 92%* 100%* 100%* 100%* 56%
4. NNR-Linear 56% 62% 98%* - 100%* 100%* 100%* 100%* 88%
5. IntelPooling 0%* 0%* 8%* 0%* - 96%* 100%* 100%* 43%
6. Neural-Linear 0%* 0%* 0%* 0%* 4%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
@@
1 2 3 4 5 6 7 8 Avg
1. DML-TS-NNR - 60%* 100%* 100%* 100%* 100%* 100%* 100%* 94%
2. DML-TS-NNR-BLM 40%* - 100%* 100%* 100%* 100%* 100%* 100%* 91%
3. DML-TS-SU 0%* 0%* - 28%* 78%* 100%* 100%* 100%* 58%
4. NNR-Linear 0%* 0%* 72%* - 86%* 100%* 100%* 100%* 65%
5. IntelPooling 0%* 0%* 22%* 14%* - 98%* 100%* 100%* 48%
6. Neural-Linear 0%* 0%* 0%* 0%* 2%* - 100%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Table 3: Pairwise comparisons between methods in the three settings of the simulation with a rectangular array of data. As in Table 2, each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The qualitative results are similar to those of Table 2. DML-TS-NNR and DML-TS-NNR-BLM perform well across all three settings and substantially outperform all other methods in the Nonlinear setting.
Table 3 displays results from pairwise method comparisons, similar to those shown in Table 2. Again the results are qualitatively similar to but slightly more exaggerated than those from the main simulation. In particular, in this simulation study, DML-TS-NNR and DML-TS-NNR-BLM outperform all other methods in 100% of repetitions in the nonlinear setting. One notable difference compared to the main simulation is that NNR-Linear and IntelPooling slightly outperform DML-TS-NNR and DML-TS-NNR-BLM in the first two settings of this simulation.
Appendix C Additional Details for Valentine Study
Personalizing treatment delivery in mobile health is a common application for online learning algorithms. We focus here on the Valentine study, a prospective, randomized-controlled, remotely administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients (Jeganathan et al., 2022; Golbus et al., 2023). We aim to use smartwatch data (Apple Watch and Fitbit) obtained from the Valentine study to learn the optimal timing of notification delivery given the users’ current context.
C.1 Data from the Valentine Study
Prior to the start of the trial, baseline data was collected from each of the participants (e.g., age, gender, baseline activity level, and health information). During the study, participants are randomized to either receive a notification ($A_{t}=1$) or not ($A_{t}=0$) at each of 4 daily time points (morning, lunchtime, mid-afternoon, evening), with probability 0.25. Contextual information was collected frequently (e.g., number of messages sent in prior week, step count variability in prior week, and pre-decision point step-counts).